OpenStudy (anonymous):

can someone help me? my problem is: -2(x-3)=8x+1

OpenStudy (anonymous):

ya sure i can help

OpenStudy (anonymous):

disturbute -2: -2x+3=8x+1 -10x=4 x=-10/4

OpenStudy (anonymous):

ok and thank you =^)

OpenStudy (anonymous):

can u give me a medal

OpenStudy (anonymous):

thank you